# MCbiF: Measuring Topological Autocorrelation in Multiscale Clusterings via 2-Parameter Persistent Homology

### v1

@@ Abstract
Datasets often possess an intrinsic multiscale structure with meaningful descriptions at different levels of coarseness. Such datasets are naturally described as multi-resolution clusterings, i.e., not necessarily hierarchical sequences of partitions across scales. To analyse and compare such sequences, we use tools from topological data analysis and define the Multiscale Clustering Bifiltration (MCbiF), a 2-parameter filtration of abstract simplicial complexes that encodes cluster intersection patterns across scales. The MCbiF can be interpreted as a higher-order extension of Sankey diagrams and reduces to a dendrogram for hierarchical sequences. We show that the multiparameter persistent homology (MPH) of the MCbiF yields a finitely presented and block decomposable module, and its stable Hilbert functions characterise the topological autocorrelation of the sequence of partitions. In particular, at dimension zero, the MPH captures violations of the refinement order of partitions, whereas at dimension one, the MPH captures higher-order inconsistencies between clusters across scales. We demonstrate through experiments the use of MCbiF Hilbert functions as topological feature maps for downstream machine learning tasks. MCbiF feature maps outperform information-based baseline features on both regression and classification tasks on synthetic sets of non-hierarchical sequences of partitions. We also show an application of MCbiF to real-world data to measure non-hierarchies in wild mice social grouping patterns across time.
Keywords: topological data analysis, multiparameter persistent homology, multiscale clustering, non-hierarchical clustering, Sankey diagrams
@@ 1 Introduction
In many applications, datasets possess an intrinsic multiscale structure, whereby meaningful descriptions exist at different scales, i.e., at different resolutions or levels of coarseness. Think, for instance, of the multi-resolution structure in commuter mobility patterns [Alessandretti et al., 2020, Schindler et al., 2023], communities in social networks [Beguerisse-Díaz et al., 2017] and thematic groups of documents [Blei et al., 2003, Grootendorst, 2022]; the subgroupings in single-cell data [Hoekzema et al., 2022] or phylogenetic trees [Chan et al., 2013]; and the functional substructures in proteins [Delvenne et al., 2010, Delmotte et al., 2011]. In such cases, the natural description of the dataset goes beyond a single clustering and consists of a multi-resolution sequence of partitions across scales parametrised by a scale parameter $t$. Traditionally, multiscale descriptions have emerged from hierarchical clustering, where $t$ corresponds to the depth of the dendrogram [Carlsson and Mémoli, 2010, Murtagh and Contreras, 2012]. However, in many important real-world applications, the data structure is multiscale, yet non-hierarchical. Examples include temporal clustering, where $t$ corresponds to physical time [Rosvall and Bergstrom, 2010, Liechti and Bonhoeffer, 2020, Bovet et al., 2022]; topic modelling and document classification, where $t$ captures the coarseness of the topic groupings [Altuncu et al., 2019, Fukuyama et al., 2023, Liu et al., 2025]; and generic multiscale clusterings for data that result from exploiting a diffusion on the data geometry, where $t$ is the increasing time horizon of the diffusion [Coifman et al., 2005, Azran and Ghahramani, 2006, Lambiotte et al., 2014].
A natural problem is then how to analyse and compare non-hierarchical multi-resolution sequences of partitions that are organised by the scale $t$. Here we address this question from the perspective of topological data analysis [Carlsson and Zomorodian, 2009, Carlsson et al., 2009, Botnan and Lesnick, 2023] by introducing the Multiscale Clustering Bifiltration (MCbiF), a 2-parameter filtration of abstract simplicial complexes that encodes the patterns of cluster intersections across all scales.
@@
A partition $\pi$ of a finite set $X=\{x_{1},x_{2},...,x_{N}\}$ is a collection of mutually exclusive subsets $C_{i}\subseteq X$ (here called clusters) that cover $X$, i.e., $\pi=\{C_{1},\dots,C_{c}\}$ such that $X=~\bigcup_{i=1}^{c}C_{i}$, and $C_{i}\bigcap C_{j}=\emptyset,\,\forall i\neq j$. The cardinality $|\pi|=c$ is the number of clusters in $\pi$ and, for notational convenience, we use $\pi_{i}$ to denote the $i$-th cluster $C_{i}$ of $\pi$.
Let $\Pi_{X}$ denote the space of partitions of $X$. We write $\pi\leq\pi^{\prime}$ if every cluster in $\pi$ is contained in a cluster of $\pi^{\prime}$. This refinement relation constitutes a partial order and leads to the partition lattice $(\Pi_{X},\leq)$ with lower bound $\hat{0}:=\{\{x_{1}\},\dots,\{x_{N}\}\}$ and upper bound $\hat{1}:=\{X\}$ [Birkhoff, 1967].
Here, we consider a sequence of partitions defined as a piecewise-constant function $\theta:[t_{1},\infty)\rightarrow\Pi_{X},\,t\mapsto\theta(t)\in\Pi_{X}$ such that a partition of $X$ is assigned to each $t$, and the scale index $t\in[t_{1},\infty)$ has $M$ change points $t_{1}<t_{2}<...<t_{M}$. In particular, $\theta(t)=\theta(t_{m})$ for $t\in[t_{m},t_{m+1})$, $m=1,\dots,M-1$, and $\theta(t)=\theta(t_{M})$ for $t\in[t_{M},\infty)$. The sequence $\theta$ is called hierarchical if either $\theta(s)\leq\theta(t),\,\forall s\leq t$, or $\theta(s)\geq\theta(t),\,\forall s\leq t$. The sequence $\theta$ is called coarse-graining if $|\theta(s)|\geq|\theta(t)|,\,\forall s\leq t$.  [^1] Coarse-graining is equivalent to non-decreasing mean cluster size (see Remark 37 in Appendix A.1). Conversely, $\theta$ is called fine-graining if $|\theta(s)|\leq|\theta(t)|,\,\forall s\leq t$.
Our goal is to characterise and analyse arbitrary sequences of partitions $\theta$, including non-hierarchical ones, in an integrated manner, taking account of memory effects across the scale $t$.
@@
Here, we are not concerned with the task of computing the multiscale clustering (i.e., the sequence of partitions $\theta$) from dataset $X$, for which several methods exist. Rather, we take $\theta$ as a given, and we aim to analyse its structure.
@@
This problem is distinct from consensus clustering, which aims to produce a summary partition by combining a set of partitions obtained, e.g., from different optimisations or clustering algorithms [Strehl and Ghosh, 2002, Vega-Pons and Ruiz-Shulcloper, 2011].
Figure 1: (a) Illustration of how the MCbiF encodes the structure of a non-hierarchical sequence of partitions $\theta$ as a bifiltration of abstract simplicial complexes $K^{s,t}$. See Example 42 for a detailed description. (b) The Hilbert functions $\mathrm{HF}_{k}(s,t)$ of the MCbiF are invariants that capture the topological autocorrelation of $\theta$: violations of the refinement order at dimension $k=0$, and higher-order cluster inconsistencies at dimension $k=1$. The Hilbert functions can be used as feature maps for downstream machine learning tasks.
@@
To address this problem, we define the MCbiF, a bifiltration of abstract simplicial complexes, which represents the clusters and their intersection patterns in the sequence $\theta$ for varying starting scale $s$ and lag tst-s (Fig. 1). Using the machinery of multiparameter persistent homology (MPH) [Carlsson and Zomorodian, 2009, Carlsson et al., 2009, Botnan and Lesnick, 2023], we prove that the MCbiF leads to a block decomposable persistence module with stable Hilbert functions $\mathrm{HF}_{k}(s,t)$, and we show that these invariants serve as measures of the topological autocorrelation of the sequence of partitions $\theta$ across the scale $t$. In particular, $\mathrm{HF}_{k}(s,t)$ quantifies the non-hierarchy in $\theta$ in two complementary ways: at dimension $k=0$, it detects the lack of a maximal partition in the subposet $\theta([s,t])$ with respect to refinement, and at dimension $k=1$, it quantifies the higher-order inconsistencies of cluster assignments across scales. In contrast, baseline methods such as ultrametrics [Carlsson and Mémoli, 2010] or information-based measures [Meilă, 2003] are restricted to pairwise comparisons between, respectively, elements or clusters; hence these methods cannot detect higher-order cluster inconsistencies. Furthermore, we provide an equivalent nerve-based construction of the MCbiF that can be interpreted as a higher-order extension of the Sankey diagram of the sequence of partitions. In the hierarchical case, the 1-dimensional MPH of the MCbiF is trivial, and the Sankey diagram reduces to a dendrogram such that the 0-dimensional MCbiF Hilbert function can be obtained from the number of branches in the dendrogram.
The Hilbert functions of the MCbiF provide interpretable feature maps that can be used in downstream machine learning tasks. In our experiments, the MCbiF feature maps outperform information-based baseline features [Meilă, 2007] on both regression and classification tasks on non-hierarchical sequences of partitions. We also show an application of MCbiF to real-world data to measure non-hierarchies in wild mice social grouping patterns across time [Bovet et al., 2022].
@@ 2 Related Work
A hierarchical, coarse-graining sequence $\theta$ with $\theta(t_{1}=0)=\hat{0}$ and $\theta(t_{M})=\hat{1}$ is called an agglomerative dendrogram, and can be represented by an acyclic rooted merge tree [Jain et al., 1999, Carlsson and Mémoli, 2010]. One can define an ultrametric $D_{\theta}$ from the first-merge times, which corresponds to the depth in the dendrogram. Carlsson and Mémoli [2010] showed that there is a one-to-one correspondence between agglomerative dendrograms and ultrametrics, which can be used to efficiently compare two such dendrograms via the Gromov-Hausdorff distance between the ultrametric spaces [Mémoli et al., 2023]. When $\theta$ is non-hierarchical, however, the first-merge times no longer define the sequence uniquely because clusters that have merged can split off again. In this case, $\theta$ cannot be represented by a tree and $D_{\theta}$ does not fulfil the triangle inequality in general. Hence, ultrametrics cannot be used to analyse and compare non-hierarchical sequences of partitions (see Section 5).
Information-based Comparison of Clusterings.
@@
Information-based measures can be used to compare a pair of partitions. Assuming a uniform distribution on $X$, one can derive probability distributions for partitions interpreted as random variables and thus measure the information gain and loss between two partitions using the conditional entropy (CE) or the variation of information (VI), which is a metric on $\Pi_{X}$ [Meilă, 2003, 2007]. See Appendix E for detailed formulas. Extending information-based measures to more than two partitions is non-trivial. In consensus clustering, the average VI is used as a consensus index (CI) [Vinh and Epps, 2009, Vinh et al., 2010] for multiple partitions. However, the CI is independent of the ordering in the sequence and so cannot capture memory effects in sequences of partitions. Another limitation of these measures is that they rely only on the joint probability between pairs of random variables, hence higher-order cluster inconsistencies are not captured (see Section 5).
3 Background
3.1 Sankey Diagrams
Non-hierarchical sequences of partitions $\theta$ are visualised by $M$-layered flow graphs $S(\theta)=(V=V_{1}\uplus...\uplus V_{M},E=E_{1}\uplus...\uplus E_{M-1})$ called Sankey diagrams [Sankey, 1898, Zarate et al., 2018], where each level $m=1,\ldots,M$ corresponds to a partition and vertices $V_{m}$ represent its clusters while the directed edges $E_{m}$ between levels indicate the overlap between clusters:
$V_{m}:=\{(m,i)\;|\;1\leq i\leq|\theta(t_{m})|\}\quad\text{and}\quad E_{m}=\left\{\left[(m,i),(m+1,j)\right]\;|\;\theta(t_{m})_{i}\cap\theta(t_{m+1})_{j}\neq\emptyset\right\},$ (1)
where $u,v]\in E_{m}$ denotes a directed edge from $u\in V_{m}$ to $v\in V_{m+1}$. If $\theta$ is hierarchical, the Sankey diagram $S(\theta)$ is a directed tree—a merge-tree if $\theta$ is coarse-graining, or a split-tree if $\theta$ is fine-graining. The graph $S(\theta)$ is sometimes also called an alluvial diagram [Rosvall and Bergstrom, 2010].
Sankey diagrams are studied in computer graphics as they allow for the visualisation of complex relational data. In this context, a Sankey diagram is represented as a layout on the plane, whereby the nodes in each layer $V_{m}$ are vertically ordered according to a ranking $\tau_{m}:V_{m}\rightarrow\{1,\dots,|V_{m}|\}$, and the layout of the Sankey diagram is then defined by the collection of such rankings, $\tau:=(\tau_{1},\dots,\tau_{M})$. For visualisation purposes, the layered layout should ideally minimise the number of crossings between consecutive layers, where a crossing between two edges $u,v],[u^{\prime},v^{\prime}]\in E_{m}$ occurs if $\tau_{m}(u)>\tau_{m}(u^{\prime})$ and $\tau_{m+1}(v)<\tau_{m+1}(v^{\prime})$ or vice versa, and the crossing number [Warfield, 1977] is given by:
$\overline{\kappa}_{\theta}:=\min_{\tau}\kappa_{\theta}(\tau).$ (3)
This problem is known to be NP-complete [Garey and Johnson, 2006] and finding efficient optimisation algorithms is an active research area [Zarate et al., 2018, Li et al., 2025].
3.2 Multiparameter Persistent Homology
Multiparameter persistent homology (MPH) is an extension of standard persistent homology to $n>1$ parameters, first introduced by Carlsson and Zomorodian [2009]. We present here basic definitions, see Carlsson and Zomorodian [2009], Carlsson et al. [2009], Botnan and Lesnick [2023] for details.
Simplicial Complex.
Let $K$ be a simplicial complex defined for the set $X$, such that $K\subseteq 2^{X}$ and $K$ is closed under the operation of building subsets. The elements of $\sigma\in K$ are called simplices and a $k$-dimensional simplex (or $k$-simplex) can be represented as $\sigma=[x_{1},...,x_{k+1}]$ where $x_{1},\dots,x_{k+1}\in X$ and we have fixed an arbitrary order on $X$. Note that $k=0$ corresponds to vertices, $k=1$ to edges, and $k=2$ to triangles. We define the $k$-skeleton $K_{k}$ of $K$ as the union of its $n$-simplices for $n\leq k$. We also define $\dim(K)$ as the largest dimension of any simplex in $K$.
Multiparameter Filtration.
Let us define the parameter space $(P,\leq)$ as the product of $n\geq 1$ partially ordered sets $P=P_{1}\times\dots\times P_{n}$, i.e., $\boldsymbol{a}\leq\boldsymbol{b}$ for $\boldsymbol{a},\boldsymbol{b}\in P$ if and only if $\boldsymbol{a}_{i}\leq\boldsymbol{b}_{i}$ in $P_{i}$ for $i=1,\dots,n$. A collection of subcomplexes $(K^{\boldsymbol{a}})_{\boldsymbol{a}\in\mathbb{R}^{n}}$ with $K=\bigcup_{\boldsymbol{a}\in\mathbb{R}^{n}}K^{\boldsymbol{a}}$ and inclusion maps $\{i_{\boldsymbol{a},\boldsymbol{b}}:K^{\boldsymbol{a}}\rightarrow K^{\boldsymbol{b}}\}_{\boldsymbol{a}\leq\boldsymbol{b}}$ that yield a commutative diagram is called a multiparameter filtration (or bifiltration for $n=2$). We denote by $\mathrm{birth}(\sigma)\subseteq P$ the set of parameters, called multigrades (or bigrades for $n=2$), at which simplex $\sigma\in K$ first appears in the filtration. For example, the sublevel filtration $K^{\boldsymbol{a}}=\{\sigma\in K\;|\;f(\sigma)\leq\boldsymbol{a}\}$ for a filtration function $f:K\rightarrow P$ maps each simplex $\sigma$ to a unique multigrade $f(\sigma)$, i.e., $|\mathrm{birth}(\sigma)|=1$. A filtration is called one-critical if it is isomorphic to a sublevel filtration, and multi-critical otherwise.
Multiparameter Persistent Homology.
Let $H_{k}$ for $k\in\{0,\dots,\dim(K)\}$ denote the $k$-dimensional homology functor with coefficients in a field [Hatcher, 2002], see Appendix C for details. Then $H_{k}$ applied to the multiparameter filtration leads to a multiparameter persistence module, i.e., a collection of vector spaces $(H_{k}(K^{\boldsymbol{a}}))_{\boldsymbol{a}\in\mathbb{R}^{n}}$, which are the homology groups whose elements are the generators of $k$-dimensional non-bounding cycles, and linear maps $\{\imath_{\boldsymbol{a},\boldsymbol{b}}:=H_{k}(i_{\boldsymbol{a},\boldsymbol{b}}):H_{k}(K^{\boldsymbol{a}})\rightarrow H_{k}(K^{\boldsymbol{b}})\}_{\boldsymbol{a}\leq\boldsymbol{b}}$ that yield a commutative diagram called multiparameter persistent homology (MPH). For dimension $k=0$, $H_{k}$ captures the number of disconnected components and for $k=1$, the number of holes. Note that, for $n=1$, we recover standard persistent homology (PH)  [Edelsbrunner et al., 2002].
Hilbert Function.
While barcodes are complete invariants of 1-parameter PH ($n=1$), the more complicated algebraic structure of MPH ($n\geq 2$) does not allow for such simple invariants in general; hence, various non-complete invariants of the MPH are used in practice. We focus on the $k$-dimensional Hilbert function [Botnan and Lesnick, 2023] defined as
$\mathrm{HF}_{k}:P\rightarrow\mathbb{N}_{0},\,\mathbf{a}\mapsto\operatorname{rank}[H_{k}(i_{\boldsymbol{a},\boldsymbol{a}})]=\dim[H_{k}(K^{\mathbf{a}})],$ (4)
which maps each filtration index $\mathbf{a}$ to the $k$-dimensional Betti number of the corresponding complex $K^{\mathbf{a}}$. We call the $k$-dimensional MPH trivial if $\mathrm{HF}_{k}=0$. The Hilbert distance is then defined as the $L_{2}$ norm on the space of Hilbert functions and can be used to compare multiparameter persistence modules.

### v2

@@ Abstract
Datasets often possess an intrinsic multiscale structure with meaningful descriptions at different levels of coarseness. Such datasets are naturally described as multi-resolution clusterings, i.e., not necessarily hierarchical sequences of partitions across scales. To analyse and compare such sequences, we use tools from topological data analysis and define the Multiscale Clustering Bifiltration (MCbiF), a 2-parameter filtration of abstract simplicial complexes that encodes cluster intersection patterns across scales. The MCbiF is a complete invariant of (non-hierarchical) sequences of partitions and can be interpreted as a higher-order extension of Sankey diagrams, which reduce to dendrograms for hierarchical sequences. We show that the multiparameter persistent homology (MPH) of the MCbiF yields a finitely presented and block decomposable module, and its stable Hilbert functions characterise the topological autocorrelation of the sequence of partitions. In particular, at dimension zero, the MPH captures violations of the refinement order of partitions, whereas at dimension one, the MPH captures higher-order inconsistencies between clusters across scales. We then demonstrate through experiments the use of MCbiF Hilbert functions as interpretable topological feature maps for downstream machine learning tasks, and show that MCbiF feature maps outperform both baseline features and representation learning methods on regression and classification tasks for non-hierarchical sequences of partitions. We also showcase an application of MCbiF to real-world data of non-hierarchical wild mice social grouping patterns across time.
@@ 1 Introduction
In many applications, datasets possess an intrinsic multiscale structure, whereby meaningful descriptions exist at different resolutions or levels of coarseness. Think, for instance, of the multi-resolution structure in commuter mobility patterns (Alessandretti et al., 2020; Schindler et al., 2023), communities in social networks (Beguerisse-Díaz et al., 2017) and thematic groups of documents (Blei et al., 2003; Grootendorst, 2022); the subgroupings in single-cell data (Hoekzema et al., 2022) or phylogenetic trees (Chan et al., 2013); and the functional substructures in proteins (Delvenne et al., 2010; Delmotte et al., 2011). In such cases, the natural description of the data goes beyond a single clustering and consists of a multiscale sequence of partitions parametrised by a scale parameter $t$. Traditionally, multiscale descriptions have emerged from hierarchical clustering, with $t$ as the depth of the dendrogram (Carlsson & Mémoli, 2010; Murtagh & Contreras, 2012). However, in many real-world applications, the data structure is multiscale yet non-hierarchical. For example in temporal clustering, where $t$ corresponds to physical time (Rosvall & Bergstrom, 2010; Liechti & Bonhoeffer, 2020; Bovet et al., 2022); in topic modelling and document classification, where $t$ captures the coarseness of the topic groupings (Altuncu et al., 2019; Fukuyama et al., 2023; Liu et al., 2025); or in clustering methods that exploit a diffusion on the data geometry, where $t$ is the increasing time horizon of the diffusion (Coifman et al., 2005; Azran & Ghahramani, 2006; Lambiotte et al., 2014).
A natural problem is then how to analyse and compare non-hierarchical, multi-resolution sequences of partitions parametrised by the scale $t$. Here we address this question from the perspective of topological data analysis (Carlsson & Zomorodian, 2009; Carlsson et al., 2009; Botnan & Lesnick, 2023) by introducing the Multiscale Clustering Bifiltration (MCbiF), a 2-parameter filtration of abstract simplicial complexes that encodes the patterns of cluster intersections across all scales.
@@
A partition $\pi$ of set $X=\{x_{1},x_{2},...,x_{N}\}$ is a collection of mutually exclusive subsets $C_{i}\subseteq X$ (or clusters) that cover $X$, i.e., $\pi=\{C_{1},\dots,C_{c}\}$ such that $X=~\bigcup_{i=1}^{c}C_{i}$ and $C_{i}\bigcap C_{j}=\emptyset,\,\forall i\neq j$. We denote the cardinality as $|\pi|=c$, and $\pi_{i}$ for the $i$-th cluster $C_{i}$ of $\pi$. Let $\Pi_{X}$ denote the space of partitions of $X$. We write $\pi\leq\pi^{\prime}$ if every cluster in $\pi$ is contained in a cluster of $\pi^{\prime}$. This refinement relation constitutes a partial order and leads to the partition lattice $(\Pi_{X},\leq)$ with lower bound $\hat{0}:=\{\{x_{1}\},\dots,\{x_{N}\}\}$ and upper bound $\hat{1}:=\{X\}$ (Birkhoff, 1967).
Here, we consider a sequence of partitions defined as $\theta:[t_{1},\infty)\rightarrow\Pi_{X},\,t\mapsto\theta(t)\in\Pi_{X}$, i.e., a piecewise-constant function that assigns a partition of $X$ to each $t$. The scale $t$ has $M$ change points $t_{1}<t_{2}<...<t_{M}$, so that $\theta(t)=\theta(t_{m})$ for $t\in[t_{m},t_{m+1})$, $m=1,\dots,M-1$, and $\theta(t)=\theta(t_{M})$ for $t\in[t_{M},\infty)$. The sequence $\theta$ is hierarchical in $[s,t]$ if we have a strict sequence of refinements: either agglomerative ($\theta(r_{1})\leq\theta(r_{2}),\forall r_{1},r_{2}\in[s,t]$ with $r_{1}\leq r_{2}$) or divisive ($\theta(r_{1})\geq\theta(r_{2}),\,\forall r_{1},r_{2}\in[s,t]$ with $r_{1}\leq r_{2}$). We say that $\theta$ is strictly hierarchical if it is hierarchical in $t_{1},\infty)$. The sequence $\theta$ is called coarse-graining if $|\theta(s)|\geq|\theta(t)|,\,\forall s\leq t$. [^1] Coarse-graining is equivalent to non-decreasing mean cluster size (see Remark 23 in Appendix B.1). Conversely, $\theta$ is called fine-graining if $|\theta(s)|\leq|\theta(t)|,\,\forall s\leq t$.
Our goal is to characterise and analyse sequences of partitions $\theta$, including non-hierarchical ones, in an integrated manner, taking account of memory effects across the scale $t$.
@@
Here, we are not concerned with the task of computing the multiscale clustering (i.e., the sequence of partitions $\theta$) from dataset $X$, for which several methods exist. Rather, we take $\theta$ as a given, and aim to analyse its structure.
@@
This problem is distinct from consensus clustering, which aims to produce a summary partition by combining a set of partitions obtained, e.g., from different optimisations or clustering algorithms (Strehl & Ghosh, 2002; Vega-Pons & Ruiz-Shulcloper, 2011).
Figure 1: (a) The MCbiF encodes the structure of a non-hierarchical sequence of partitions $\theta$ as a bifiltration of abstract simplicial complexes $K^{s,t}$ (see Example 37). (b) The Hilbert functions $\mathrm{HF}_{k}(s,t)$ of the MCbiF are invariants that capture the topological autocorrelation of $\theta$: violations of the refinement order at dimension $k=0$, and higher-order cluster inconsistencies at dimension $k=1$. The $\mathrm{HF}_{k}(s,t)$ can be used as feature maps for downstream machine learning tasks.
@@
To address this problem, we define the MCbiF, a bifiltration of abstract simplicial complexes, which represents the clusters and their intersection patterns in the sequence $\theta$ for varying starting scale $s$ and lag tst-s (Fig. 1). The MCbiF is a complete invariant of $\theta$, and we use the machinery of multiparameter persistent homology (MPH) (Carlsson & Zomorodian, 2009; Carlsson et al., 2009; Botnan & Lesnick, 2023) to summarise its topological structure. We prove that the MCbiF leads to a block decomposable persistence module with stable Hilbert functions $\mathrm{HF}_{k}(s,t)$, and we show that these invariants characterise the topological autocorrelation of the sequence of partitions $\theta$ across scales. In particular, the $\mathrm{HF}_{k}(s,t)$ quantify the non-hierarchy in $\theta$ in complementary ways: at dimension $k=0$, it detects the lack of a maximal partition in the subposet $\theta([s,t])$ with respect to refinement; at dimension $k=1$, it quantifies higher-order inconsistencies of cluster assignments across scales. In contrast, baseline methods such as ultrametrics (Carlsson & Mémoli, 2010) or information-based measures (Meilă, 2003) are restricted to pairwise comparisons between elements or clusters, respectively; hence, they cannot detect higher-order cluster inconsistencies. Furthermore, we provide an equivalent nerve-based construction of the MCbiF that can be interpreted as a higher-order extension of the Sankey diagram of the sequence of partitions. In particular, for the hierarchical case, the 0-dimensional MCbiF Hilbert function can be obtained from the number of branches in the Sankey diagram, which reduces to a dendrogram. Finally, we show that $\mathrm{HF}_{k}(s,t)$ provide interpretable feature maps usable for downstream machine learning tasks. In our experiments, the MCbiF feature maps outperform both baseline features and representation learning methods on regression and classification tasks for non-hierarchical sequences of partitions. We also showcase an application of MCbiF to real-world data of non-hierarchical wild mice social grouping patterns across time (Bovet et al., 2022).
@@ 2 Related Work
A hierarchical, coarse-graining sequence $\theta$ with $\theta(t_{1}=0)=\hat{0}$ and $\theta(t_{M})=\hat{1}$ is called an agglomerative dendrogram, and can be represented by an acyclic rooted merge tree (Jain et al., 1999; Carlsson & Mémoli, 2010). One can define an ultrametric $D_{\theta}$ from the first-merge times, which follows from the depth in the dendrogram. Carlsson & Mémoli (2010) showed there is a one-to-one correspondence between agglomerative dendrograms and ultrametrics that can be used to efficiently compare dendrograms using the Gromov-Hausdorff distance between the ultrametric spaces (Mémoli et al., 2023). When $\theta$ is non-hierarchical, however, first-merge times no longer define the sequence uniquely, as merged clusters can split again. Hence, $\theta$ cannot be represented by a tree, and $D_{\theta}$ does not fulfil the triangle inequality in general. As a result, ultrametrics cannot be used to analyse and compare non-hierarchical sequences of partitions (see Section 4).
Pairwise Comparison of Partitions.
@@
Different measures to compare pairs of partitions have been introduced in the literature. The Adjusted Rand Index (ARI) is the chance-corrected Rand Index that compares clusterings by counting elements that are assigned to the same or different clusters (Hubert & Arabie, 1985). Information-based measures compute the information gain and loss between two partitions using the Conditional Entropy (CE) or the Variation of Information (VI), which is a metric on $\Pi_{X}$ (Meilă, 2003; 2007) (see formulas in Appendix F). The Maximum Overlap Distance (MOD), which is also a metric on $\Pi_{X}$, measures the minimal classification error when one partition is assumed to be the correct one  (Peixoto, 2021). A key limitation of all these measures is that they rely only on pairs of partitions and cannot capture higher-order cluster inconsistencies; yet extending them to more than two partitions is non-trivial. In consensus clustering, the average of the pairwise ARI, VI or MOD between the partitions in a set is used as a consensus index (Vinh & Epps, 2009; Vinh et al., 2010). However, these average measures are insensitive to ordering and so cannot capture memory effects in sequences of partitions.